What is the standard form of a parabola with vertex at (-5,1) and focus at (2,1) ?
Answer by lwsshak3(11628)

What is the standard form of a parabola with vertex at (-5,1) and focus at (2,1) ?
**
Standard form of equation for a parabola which opens rightward: (y-k)^2=4p(x-h), with (h,k) being the (x,y) coordinates of the vertex.
For given parabola:
vertex: (-5,1) (given)
axis of symmetry: y=1
p=7 (distance from the vertex to the focus on the axis of symmetry)
4p=28
Equation:
(y-1)^2=28(x+5)